# BST: Badminton Stroke-type Transformer for Skeleton-based Action Recognition in Racket Sports

### v2

@@ B Training Settings
The implementation is based on the PyTorch framework. All downstream classification models in Tab. 1 were trained from scratch using cross-entropy loss. Moreover, the compared models are derived from their official public implementations for better code readability and more efficient training and inference. We chose AdamW with default parameters as the optimizer and cosine annealing with warm-up learning rate scheduler to adjust the learning rate during training. The hyperparameters used in our experiments are listed below:
@@
  • learning_rate: $5\times 10^{-4}$
@@
  • cosine_annealing_num_cycles: $0.25$
  • n_classes: $35$
[…]
C Classes in the Dataset
There are $35$ classes in the dataset as shown in Tab. A.
Class ID Num of Strokes Player 中文類別名稱 Class Name
0 2982 Top 放小球 net shot
1 1618 Top 擋小球 return net
2 1287 Top 殺球 smash
3 831 Top 點扣 wrist smash
4 2343 Top 挑球 lob
5 123 Top 防守回挑 defensive return lob
6 1309 Top 長球 clear
7 319 Top 平球 drive
8 234 Top 後場抽平球 back-court drive
9 1036 Top 切球 drop
10 615 Top 過渡切球 passive drop
11 1317 Top 推球 push
12 270 Top 撲球 rush
13 144 Top 防守回抽 defensive return drive
14 606 Top 勾球 cross-court net shot
15 904 Top 發短球 short service
16 159 Top 發長球 long service
17 2842 Bottom 放小球 net shot
18 1756 Bottom 擋小球 return net
19 1075 Bottom 殺球 smash
20 728 Bottom 點扣 wrist smash
21 2536 Bottom 挑球 lob
22 155 Bottom 防守回挑 defensive return lob
23 1352 Bottom 長球 clear
24 335 Bottom 平球 drive
25 201 Bottom 後場抽平球 back-court drive
26 943 Bottom 切球 drop
27 583 Bottom 過渡切球 passive drop
28 1335 Bottom 推球 push
29 201 Bottom 撲球 rush
30 238 Bottom 防守回抽 defensive return drive
31 620 Bottom 勾球 cross-court net shot
32 954 Bottom 發短球 short service
33 200 Bottom 發長球 long service
34 1278 - 未知球種 none
Table A: Classes in the dataset.

### v3

@@ Appendix D Training Settings
The implementation is based on the PyTorch framework. All downstream classification models in this paper were trained from scratch using cross-entropy loss. Moreover, the compared models are derived from their official public implementations. During training, we applied random shift data augmentation with a probability of $0.3$, where the shift value was uniformly sampled from [0.3,0.3)[-0.3,0.3). We chose AdamW as the optimizer and cosine annealing with warm-up learning rate scheduler to adjust the learning rate. We validated the model on the validation set after each epoch and saved the model with the best macro-F1 score on the validation set. The hyperparameters used in our experiments are listed below:
@@
  • learning_rate:
    • $5\times 10^{-2}$ on BlockGCN
    • $1\times 10^{-3}$ on ST-GCN, ProtoGCN, SkateFormer
    • $5\times 10^{-4}$ on other models
  • cosine_annealing_num_cycles: $0.25$
@@
  • weight_decay:
    • $5\times 10^{-4}$ on SkateFormer partial weights
    • $1\times 10^{-2}$ on other models
  • label_smoothing: $0.1$
  • sequence_length:
    • $30$ on ShuttleSet with fixed-width strategy
    • $100$ on ShuttleSet with our strategy
    • $72$ on BadmintonDB
    • $100$ on TenniSet
  • n_classes:
    • $35$ on ShuttleSet (excessive granularity)
    • $25$ on ShuttleSet (merged)
    • $18$ on BadmintonDB
    • 66 on TenniSet
Appendix E Details of the Datasets
ShuttleSet (excessive granularity).   The original dataset consists of 19 distinct stroke categories, including a ”none” type. One category, driven flight, containing only 52 strokes across the entire dataset was merged into the ”none” category by us.
Merged types:
   $\scriptstyle\bullet$ ”driven flight” $\Rightarrow$ ”none”
Additionally, since the dataset does not differentiate between strokes played by the top or bottom player within each category (except for ”none”), the number of stroke categories (except for ”none”) was doubled. This adjustment resulted in a total of 35 categories, including ”none”, as shown in Tab. F.
The court information is provided in the dataset, so we don’t need to detect it by ourselves. If there were less than two people in the court, we cleared the information (poses and shuttlecock trajectory) of that frame to zero, since that frame was definitely not in a standard camera perspective.
Class ID #Train #Val #Test #Total Player 中文類別名稱 Class Name
0 2402 357 223 2982 Top 放小球 net shot
1 1257 179 182 1618 Top 擋小球 return net
2 980 185 122 1287 Top 殺球 smash
3 605 120 106 831 Top 點扣 wrist smash
4 1858 266 219 2343 Top 挑球 lob
5 96 15 12 123 Top 防守回挑 defensive return lob
6 925 197 187 1309 Top 長球 clear
7 258 28 33 319 Top 平球 drive
8 172 27 35 234 Top 後場抽平球 back-court drive
9 782 153 101 1036 Top 切球 drop
10 463 94 58 615 Top 過渡切球 passive drop
11 1005 156 156 1317 Top 推球 push
12 203 33 34 270 Top 撲球 rush
13 96 26 22 144 Top 防守回抽 defensive return drive
14 503 60 43 606 Top 勾球 cross-court net shot
15 733 84 87 904 Top 發短球 short service
16 97 19 43 159 Top 發長球 long service
17 2282 338 222 2842 Bottom 放小球 net shot
18 1381 207 168 1756 Bottom 擋小球 return net
19 780 152 143 1075 Bottom 殺球 smash
20 589 58 81 728 Bottom 點扣 wrist smash
21 1974 347 215 2536 Bottom 挑球 lob
22 109 32 14 155 Bottom 防守回挑 defensive return lob
23 946 214 192 1352 Bottom 長球 clear
24 263 30 42 335 Bottom 平球 drive
25 132 36 33 201 Bottom 後場抽平球 back-court drive
26 741 112 90 943 Bottom 切球 drop
27 452 65 66 583 Bottom 過渡切球 passive drop
28 1030 169 136 1335 Bottom 推球 push
29 155 31 15 201 Bottom 撲球 rush
30 170 32 36 238 Bottom 防守回抽 defensive return drive
31 498 61 61 620 Bottom 勾球 cross-court net shot
32 763 90 101 954 Bottom 發短球 short service
33 114 27 59 200 Bottom 發長球 long service
34 927 241 162 1330 - 未知球種 none
#Sum 25741 4241 3499 33481 - - -
Table F: Classes in ShuttleSet (excessive granularity).
ShuttleSet (merged).   Everything is the same as the excessive granularity version, except for the following merged categories:
   $\scriptstyle\bullet$ ”wrist smash” $\Rightarrow$ ”smash”
   $\scriptstyle\bullet$ ”defensive return lob” $\Rightarrow$ ”lob”
   $\scriptstyle\bullet$ ”driven flight” $\Rightarrow$ ”drive”
   $\scriptstyle\bullet$ ”back-court drive” $\Rightarrow$ ”drive”
   $\scriptstyle\bullet$ ”defensive return drive” $\Rightarrow$ ”drive”
   $\scriptstyle\bullet$ ”passive drop” $\Rightarrow$ ”drop”
Thus, there are 25 classes in this dataset as shown in Tab. G.
Class ID #Train #Val #Test #Total Player 中文類別名稱 Class Name
0 875 241 162 1278 - 未知球種 none
1 2402 357 223 2982 Top 放小球 net shot
2 1257 179 182 1618 Top 擋小球 return net
3 1585 305 228 2118 Top 殺球 smash
4 1954 281 231 2466 Top 挑球 lob
5 925 197 187 1309 Top 長球 clear
6 548 81 90 719 Top 平球 drive
7 1245 247 159 1651 Top 切球 drop
8 1005 156 156 1317 Top 推球 push
9 203 33 34 270 Top 撲球 rush
10 503 60 43 606 Top 勾球 cross-court net shot
11 733 84 87 904 Top 發短球 short service
12 97 19 43 159 Top 發長球 long service
13 2282 338 222 2842 Bottom 放小球 net shot
14 1381 207 168 1756 Bottom 擋小球 return net
15 1369 210 224 1803 Bottom 殺球 smash
16 2083 379 229 2691 Bottom 挑球 lob
17 946 214 192 1352 Bottom 長球 clear
18 595 98 111 804 Bottom 平球 drive
19 1193 177 156 1526 Bottom 切球 drop
20 1030 169 136 1335 Bottom 推球 push
21 155 31 15 201 Bottom 撲球 rush
22 498 61 61 620 Bottom 勾球 cross-court net shot
23 763 90 101 954 Bottom 發短球 short service
24 114 27 59 200 Bottom 發長球 long service
#Sum 25741 4241 3499 33481 - - -
Table G: Classes in ShuttleSet (merged).
BadmintonDB.   Merged types:
   $\scriptstyle\bullet$ ”Flick-Serve” $\Rightarrow$ ”Serve”
Thus, there are 18 classes in this dataset as shown in Tab. H.
In court detection (using MonoTrack [25]), since there is only one standard camera perspective in a badminton match, it only needs to be performed once per match video, on a single frame that meets this criterion. If the court detection fails on that frame, actually, it will likely fail on all frames of that match video. We marked the four corners of the court manually in this case. If there were less than two people in the court, we conducted the same operation as in ShuttleSet.
We used duplicated data to balance the classes during training.
The reason for SkateFormer [10] not in the comparison table is that setting a proper sequence length for it is not an easy task.
Class ID #Train #Val #Test #Total Player Class Name
0 497 62 63 622 Bottom Block
1 263 32 34 329 Bottom Clear
2 60 7 9 76 Bottom Drive
3 260 32 33 325 Bottom Dropshot
4 28 3 5 36 Bottom Net-Kill
5 734 91 93 918 Bottom Net-Lift
6 663 82 84 829 Bottom Net-Shot
7 179 22 23 224 Bottom Serve
8 392 49 49 490 Bottom Smash
9 458 57 58 573 Top Block
10 203 25 26 254 Top Clear
11 74 9 10 93 Top Drive
12 289 36 37 362 Top Dropshot
13 52 6 8 66 Top Net-Kill
14 712 89 89 890 Top Net-Lift
15 639 79 81 799 Top Net-Shot
16 157 19 21 197 Top Serve
17 460 57 58 575 Top Smash
#Sum 6120 757 781 7658 -
Table H: Classes in BadmintonDB.
TenniSet.   This dataset contains 6 classes as shown in Tab. I.
In court detection (using TennisCourtDetector [36]), it was performed on each frame, as the camera perspective often shifts slightly left or right.
When determining who the two players are among the people in a frame, since the two tennis players are often outside the court lines, instead of the situations in badminton, we choose the person(s) closest to the top or bottom side of the court line(s), if there was less than two people inside the court lines. If there were less than two people in the set we picked, we cleared the information (poses and ball trajectory) of that frame to zero, since the court detection was probably failed in this case (not in a standard camera perspective or model’s failure).
In ball trajectory tracking, we used TrackNetV3 [1] (with attention), which was originally trained on badminton videos, for convenience. Although the results were not perfect, often missing the ball during bounces, the downstream model was still able to learn useful information.
The reason for other three models not in the comparison table is that they take too long time and too many resources to train on a long sequence length (e.g. 100).
Class ID #Train #Val #Test #Total Class Name
0 432 33 135 600 HFL (Hit Far Left)
1 474 37 145 656 HFR (Hit Far Right)
2 514 37 119 670 HNL (Hit Near Left)
3 448 33 144 625 HNR (Hit Near Right)
4 412 30 90 532 SF (Serve Far)
5 339 27 120 486 SN (Serve Near)
#Sum 2619 197 753 3569 -
Table I: Classes in TenniSet. ”Far” means ”Top” and ”Near” means ”Bottom”. ”Left” means that the ball was on the left side of the player who hit it in the camera perspective (not the player’s perspective). ”Right” means the opposite.